A car in stop-and-go traffic starts at rest, moves
Answer #1

we know,

the area under the velocity time graph shows the distance moved,

the total distance travelled = 21m = area under the graph

area under the graph = area of the trapezium = 1/2 * (sum of parallel side) * height

= 1/2 (2+8) * V = 5V

hence 5V = 21

hence V= 4.2 m/s

for PART A

the distance cover = area of first triangle = 1/2 * 4 * 4.2 =8.4 m

PART B

the distance cover = area of second triangle = 1/2 * 2 * 4.2 =4.2 m

PART C

already calculated V = 4.2 m/s
Answer #2

a)3

b)6

c)2

source: mastering physics
answered by: anonymous